OpenStudy (anonymous):

martha places $200 in a checking account that will earn 3% interest compounded annually.if she close the account after two years and incurs a closing charge of $2.75, how much will she have?

OpenStudy (anonymous):

A = 200 ( 1.03) ^2 = $212.18 => 212.18 - 2.75 = $209.43